OpenStudy (anonymous):

URGENT!!! QUICK HELP! Which of these ordered pairs is a solution to the linear inequality y – 4x ≤ –6? A. (–2, 4) B. (1, –2) C. (1, 3) D. (2, 3)

OpenStudy (anonymous):

Substitute the x & y values into the problem and see if you get a true statement. (-2,4). 4-4(-2) less than or = to -6 4+8 isn't less than or = to -6

OpenStudy (anonymous):

Do that for each pair

OpenStudy (anonymous):

Can you tell me the answer i dont get it?

OpenStudy (anonymous):

Try B. (1,-2) -2 - 4(1) LT/ to -6 -2 - 4 = -6. This meets the condition of less than or equal to. B is the solution.